# Transformational Creativity in Science: A Graphical Theory

### v1

@@ Introduction
In Phase I: Normal Science of Kuhn’s model, a paradigm’s structure and constraints are assumed as a foundation to explore within the conceptual space they induce (?). Likewise, in our definition, axioms do not depend on other constraints but are instead assumed for the possibility of inquiry within the space. Also note that no connectedness assumption about the graph is made, meaning that the graph can have multiple connected components to reflect disconnected but related subgraphs of knowledge, corresponding to subsets of constraints over possible artifacts. A new artifact is generated from a conceptual space in the following way:
@@
Transformational creativity occurs when modifications are made to a conceptual space. Our definition of a space $S=(V,E)$ naturally supports transformations in two ways:
@@
Conclusion and Future Work
In this paper, we provided a graphical theory of transformational scientific creativity, illustrated the transformative potential of modifying axioms, and used our framework to elucidate the transformativeness of several groundbreaking historical discoveries. One implication of the rigid structure introduced above is that our theory may be best suited to areas of science involving formal systems, such as mathematics, engineering theory, and computer science. Nevertheless, in exchange for generality, we obtain more specificity as to how the transformational creativity process can be formulated in these domains, opening the door for future theoretical or computational inquiry. Our work, like ? (?), is a crucial first step towards formalizing transformational creativity.

### v2

@@ Introduction
In Phase I: Normal Science of Kuhn’s model, a paradigm’s structure and constraints are assumed as a foundation to explore within the conceptual space they induce (?). Likewise, in our definition, axioms do not depend on other constraints but are instead assumed for the possibility of inquiry within the space. Also note that no connectedness assumption about the graph is made, meaning that the graph can have multiple connected components to reflect disconnected but related subgraphs of knowledge, corresponding to subsets of constraints over possible artifacts. A new artifact is generated from a conceptual space in the following way.
@@
Transformational creativity occurs when modifications are made to a conceptual space. Our definition of a space $S=(V,E)$ naturally supports transformations in two ways.
@@
Towards AI Super-Scientists: Using Our Framework for Transformative Discovery
In light of a series of recent works using AI for scientific ideation (????), experiment design (?), and even end-to-end discovery (??), we believe our theory opens a promising avenue towards building the first AI super-scientists poised to make transformative breakthroughs. By coupling the wide-reaching creative abilities of large language models (LLMs) with our graphical theory above, we envision a neuro-symbolic approach to transformative scientific discovery that, given a particular scientific field or sub-field $\mathcal{D}$, can be built using the following components.
  1. 1.
    Background Knowledge: A collection $\mathcal{C}$ of scientific literature (surveys, conference proceedings, journal papers, etc.) representing background knowledge about $\mathcal{D}$.
  2. 2.
    Graph Construction: A DAG construction technique $M:\mathcal{C}\to\mathcal{G}$, where $\mathcal{G}$ is a space of graphs structured according to Definition 1, representing the dependency structure of rules and axioms in $\mathcal{D}$.
  3. 3.
    Problem Identification: The identification of a set $\mathcal{P}$ of “open problems” or anomalies that currently plague the research community in $\mathcal{D}$, obtained from the background literature $A:\mathcal{C}\to\mathcal{P}$.
  4. 4.
    Idea Generation: Finally, an idea generation module
    $I:\mathcal{L}\times\mathcal{G}\times\mathcal{P}\to\mathcal{T}\times\mathcal{G}$
    that maps from a space of large language models $\mathcal{L}$, dependency graphs $\mathcal{G}$, and open problems $\mathcal{P}$ to produce a transformative idea text $t\in\mathcal{T}$ and its corresponding graph transformation $G^{\prime}\in\mathcal{G}$.
This approach offers a natural way to incorporate knowledge graphs to enhance the use of LLMs for scientific discovery (??), a direction which has shown early promise in improving the technical detail of generated ideas (?). Moreover, by making graph transformations explicit, this technique also boasts significant interpretability benefits.
Conclusion
In this paper, we provided a graphical theory of transformational scientific creativity, illustrated the transformative potential of modifying axioms, and used our framework to elucidate the transformativeness of several groundbreaking historical discoveries.
Subsequent investigations could extend our theory by providing greater depth, breadth, and nuance to the framework, contextualized within the history and philosophy of science. For example, we suggest exploring in-depth historical illustrations for one to three exemplary discoveries, such as those in Table 1, showing how all components of the theory are instantiated and transformed. Similarly, it may be insightful to study a large collection of scientific discoveries, such as those addressed in ? (?) and ? (?), and delineate a “logic of discovery” via common graph transformation patterns among these cases. Historical studies such as these may suggest criteria for “good” graph transformations and highlight distinctions between edge and vertex modifications.
One notable implication of our theory’s rigid structure is that it may be best suited to scientific areas that involve formal systems, such as mathematics, engineering theory, and computer science. Nevertheless, in exchange for generality, we obtain more specificity regarding how the transformational creativity process can be formulated in these domains, opening the door for future theoretical or computational inquiry. Our work, like ? (?), is a crucial first step towards formalizing transformational creativity.
Acknowledgments
The authors thank Jonathan Livengood for helpful feedback on an earlier version of this paper.